OpenStudy (anonymous):

T(n)=1/(1-3n) would be expanded as 1+3n+3^2n^2+..........+3^rn^r....... how it is possible???

OpenStudy (anonymous):

1/(1-3n)=(1-3n)^-1 =1+(-1)(-3n)+(-1)^2(3n)^2+(-1)^3(-3n)^3+......+(-1)^r(-3n)^r+.... =1+3n+3^2n^2+......3^rn^r+...... this is the binomial theorem...do u know that?

OpenStudy (anonymous):

thanks......